Two equal positive charges of magnitude q1 = q2 = 4 nC are on the y axis at y1 = +3 cm and y2 = -3 cm.
(a) What is the magnitude and direction of the electric field on the x axis at x = 3 cm?
N/C + N/C

(b) What is the force exerted on a third charge q0 = 6 nC when it is placed on the x axis at x = 3 cm?
N + N
1
Expert's answer
2012-03-20T11:03:35-0400
a) E=2*q*3*sqrt(2)/4*pi*eо*r^2=2*4*10^(-9)*3*1.41/(4*3.14*8.85*10^-12*9)=34.
Direction: positive x-axis
b) F=q0*2*q*3*sqrt(2)/4*pi*eо=5*34=170N